OpenStudy (anonymous):

The function f(x) is shown below. f(x) = (5)^(x - 1) The function f(x) is shifted to the right by 13 units. Which of the following best represents the new function? f(x) = (5)^(x + 13) f(x) = (5)^(x + 12) f(x) = (5)^(x - 14) f(x) = (5)^(x - 13)

OpenStudy (sidsiddhartha):

hint: to shift f(x) to the right by 13 units, you have to replace x with x+13. So you need to evaluate f(x+13)

OpenStudy (anonymous):

So the answer is C?

OpenStudy (sidsiddhartha):

f(x) = (5)^(x - 1) so f(x+13)=(5)^(x+13 - 1)=(5)^(x+12) it's B :)

OpenStudy (anonymous):

Thanks :)

OpenStudy (anonymous):

isn't it negative when you shift to the right? Then the answer has to be C instead

OpenStudy (sidsiddhartha):

when ur shifting to the left then it would be negetive @alias

OpenStudy (primeralph):

@RosieF You were right initially. It's C.

OpenStudy (primeralph):

Left shift is positive. Right shift is negative.

OpenStudy (anonymous):

@sidsiddhartha I'm pretty sure I'm correct. But either way, if you're still confused, just come up with an example, then test a few points to see for yourself